# Forecasting and Manipulating the Forecasts of Others

### 12 Mar 2026 v1

@@ Abstract
In strategic environments with private information, evaluating a change in policy requires predicting how the equilibrium responds—but when actions reshape opponents’ signals, each agent’s optimal response depends on an infinite hierarchy of beliefs about beliefs [28] that has resisted exact analysis for four decades. We provide the first exact equilibrium characterization of finite-player continuous-time LQG games with endogenous signals. Conditioning on primitive Brownian shocks rather than the physical state—a dynamic analogue of Harsanyi’s common-prior construction—collapses the belief hierarchy onto deterministic two-time kernels, reducing Nash equilibrium to a deterministic fixed point with no truncation and no large-population limit. The characterization yields an explicit information wedge $\mathcal{V}^{i}_{t}$—a deterministic Volterra process—that prices the marginal value of shifting opponents’ posteriors. The wedge vanishes precisely when signals are exogenous to controls, formally delineating the boundary where strategic belief manipulation matters, and provides a closed-form mapping from information primitives to equilibrium outcomes.
@@ 1 Introduction
This joint determination is the central obstacle. Consider a central bank shifting from quarterly to real-time disclosure of its output gap estimates. Each firm’s pricing decision depends on its belief about the gap and on its belief about how other firms filter the bank’s announcements. Under the new policy, signal precision changes, which changes how firms filter, which changes how each firm forecasts the pricing decisions of others, which changes equilibrium inflation dynamics. Existing tools cannot trace this chain because the Townsend hierarchy makes the mapping from signal precision to equilibrium outcomes intractable. The characterization in this paper makes it explicit.
The same obstruction arises across a wide range of fields: in market microstructure, where prices aggregate dispersed information and trading itself is informative; in dynamic information design, where the mapping from signal structures to equilibrium outcomes is intractable when filtering and equilibrium interact; in contract theory, where multiple agents exert unobservable effort on a shared outcome and draw inferences from it; and in decentralized control—autonomous vehicle fleets, multi-robot systems with misaligned operators—wherever agents observe a shared state through private channels and actions reshape what others sense. Existing approaches to the decentralized control problem typically assume cooperative teams or a common information base; the framework here requires neither.
@@
The conceptual move parallels Harsanyi’s [10] resolution of the belief hierarchy in incomplete-information games. Harsanyi replaced an infinite regress of beliefs about beliefs with a common prior over a type space from which all higher-order beliefs follow by conditioning. We replace the Townsend regress with the common probability space of primitive shocks $W=(W^{0},\ldots,W^{n})$, from which all higher-order conditional expectations follow by orthogonal projection onto deterministic two-time kernels with no truncation. The analogy is imperfect in a revealing way: Harsanyi’s types are drawn once and the information partition is fixed, whereas here noise unfolds continuously and actions reshape signals, so the closure requires not just a common prior but a fixed point—the deterministic kernels governing how projections evolve must be consistent with the strategy profile they induce (Figure 1).
@@
The main results establish that the best response to any profile of deterministic-kernel strategies is itself deterministic-kernel (closure over the full admissible $L^{2}$ class), so Nash equilibrium reduces to a deterministic fixed point in two-time kernels. The equilibrium characterization produces an explicit information wedge $\mathcal{V}^{i}_{t}$: a Volterra process with deterministic mean $\bar{\mathcal{V}}^{i}(t)$ and deterministic kernel $\mathcal{V}^{i}(t,r)$, which prices the marginal value to player $i$ of shifting opponents’ posteriors. The mean wedge distorts equilibrium mean actions; the kernel wedge distorts equilibrium impulse responses. Both vanish identically when signals are exogenous to controls, formally delineating the boundary where strategic belief manipulation matters. The propagation $\Delta P^{i}\to\Delta\widetilde{X}^{i}\to\Delta\mathcal{V}^{i}_{t}\to\Delta D^{i}$ provides an explicit mapping from information primitives to equilibrium outcomes that a model with fixed filtering cannot deliver: because information enters through precision paths $\{P^{i}(\cdot)\}$ and affects the equilibrium only through deterministic filtering sensitivities, a principal can trace how disclosure precision propagates into equilibrium payoffs without simulating higher-order beliefs.
(i) Candidate Opponents use Volterra controls (ii) Closure State & filters have det. kernels (iii) Deviations Fixed maps $\Rightarrow$ det. propagation (iv) Best response FOC $\Rightarrow$ control linear in $\widehat{W}^{\,i}$ (v) Fixed point BR map closes $\Longleftrightarrow$ Nash eq.
Figure 1: The Volterra fixed-point loop. (i) Assume opponents use Volterra controls with deterministic kernels. (ii) The state and filtering equations close on deterministic two-time kernels. (iii) Under fixed strategy maps, unobserved deviations propagate deterministically. (iv) The maximum principle produces a best response that is itself Volterra, closing the loop. (v) A fixed point of this best-response map is a Nash equilibrium.
@@ 1.1 Related Literature
The decentralized decision problem with shared objectives originates with Marschak and Radner [19]; Radner [23] established certainty equivalence and linear optimal policies for LQG teams. Team-theoretic solutions bypass the belief hierarchy because aligned objectives eliminate the incentive to manipulate others’ forecasts. The information wedge $\mathcal{V}^{i}_{t}$ is precisely the object that vanishes under team objectives and survives under misaligned payoffs.
@@
Nimark [22] approximates the higher-order belief hierarchy by truncation at finite order—the closest existing method in spirit to the present paper. The deterministic kernel system derived here provides an exact benchmark against which truncation error can be assessed.
@@ 2 The Decentralized LQG Game
The class of games we study has three features: (i) multiple agents share a common evolving state but observe it through private noisy channels; (ii) each agent’s action feeds back into the state dynamics and thereby into every other agent’s signal process; and (iii) agents minimize individual quadratic objectives over a finite horizon. Because actions enter the state, they simultaneously serve the physical objective and reshape what opponents observe, making each agent’s problem jointly strategic and informational. Features (i)–(ii) generate the infinite belief hierarchy identified by Townsend; feature (iii) provides the linear–Gaussian structure that makes it tractable.
@@ 2.1 Primitives
Stacked noise and block selectors.
@@
Unresolved state kernel.
@@
$\widetilde{X}^{i}(t,u):=X(t,u)-\widehat{X}^{\,i}(t,u),$
the component of the state’s impulse response to the $u$-th shock that player $i$ has not yet resolved at time tt—large when the shock is recent or the observation channel weak, and shrinking as observations accumulate. Formally, $\widetilde{X}^{i}(t,u)=\operatorname{Cov}(X_{t},\,dW_{u}\mid\mathcal{F}_{t}^{i})/du$; it serves as the filtering gain in the noise-state update (Theorem 3.1), with its algebraic representation in terms of $F^{i}$ given in (A.12).
@@
Each player $i$ incurs the random cost
$\begin{split}\mathcal{C}^{i}(D^{i};D^{-i}):=&\int_{0}^{T}\left(X_{t}^{\top}G^{i}_{XX}(t)X_{t}+2(G^{i}_{X}(t))^{\top}X_{t}+G^{i}(t)+(D^{i}_{t})^{\top}G^{i}_{DD}(t)D^{i}_{t}\right)dt\\ &+X_{T}^{\top}G^{i}_{XX}(T)X_{T}+2G^{i}_{X}(T)X_{T},\end{split}$ (2.3)
where $G^{i}_{XX}(t)$ and $G^{i}_{DD}(t)$ are symmetric positive semidefinite matrices (with $G^{i}_{DD}(t)$ positive definite), and $D^{-i}:=(D^{j})_{j\neq i}$ denotes opponents’ controls. Each player $i$ seeks to minimize
@@ 2.2 Nash Equilibrium
In a unilateral deviation by $i$, opponents’ strategy maps $Y^{k}_{\cdot\wedge t}\mapsto D_{t}^{k}[Y^{k}]$ are held fixed and applied to the perturbed observation histories; in the Volterra class these maps are parameterized by deterministic kernels that remain unchanged off equilibrium.
@@
This paper aims to establish that the Volterra class is closed under best responses and to derive a deterministic kernel system whose fixed points correspond to Nash equilibria in that class.
@@
Definition 2.2 (Primitive-noise Volterra process).
An $L^{2}$ stochastic process $L=(L_{t})_{t\in[0,T]}$ is a (primitive-noise) Volterra process if there exist deterministic functions
@@
Remark 2.2 (Properties of primitive-noise Volterra processes).
By construction, primitive-noise Volterra processes are $\mathbb{F}$-adapted Gaussian processes. Moreover, by Itô isometry the square-integrability condition above is equivalent to $L_{t}\in L^{2}$ for each $t$. Finally, the martingale representation theorem implies that any $L^{2}$ $\mathcal{F}_{t}$-measurable random variable admits a representation of the form $\int_{0}^{t}Z_{s}\,dW_{s}$ with a stochastic integrand $Z$; the defining restriction of the Volterra class is that the representation kernel $L(t,\cdot)$ is deterministic (for each fixed $t$).
Definition 2.3 (Noise-state Volterra control / strategy).
Fix a player $i$ and recall their noise-state path $\widehat{W}_{t}^{i}(\cdot)$ (Section 3). An admissible control $D^{i}=(D_{t}^{i})_{t\in[0,T]}$ is a noise-state Volterra control if there exist deterministic functions
@@
$D_{t}^{i}=\bar{D}^{i}(t)+\int_{0}^{t}D^{i}(t,u)\,d_{u}\widehat{W}_{t}^{i}(u),\qquad\int_{0}^{t}\|D^{i}(t,u)\|^{2}\,du<\infty,$
and we extend $D^{i}(t,u)$ by 0 for $u>t$. Equivalently, writing $\widehat{W}^{\,i}[Y^{i}]$ to emphasize dependence on the observation path, this defines the strategy map
$Y^{i}_{\cdot\wedge t}\longmapsto D_{t}^{i}[Y^{i}]:=\bar{D}^{i}(t)+\int_{0}^{t}D^{i}(t,u)\,d_{u}\widehat{W}_{t}^{\,i}[Y^{i}](u).$
@@ 3 Main results
We work in the linear–Gaussian model of Section 2 and establish two closure results—filtering and best response—that reduce Nash equilibrium to a deterministic fixed point in two-time kernels. The filtering closure (Theorem 3.1) shows that noise-state dynamics have deterministic kernels; the best-response closure (Theorem 3.2) derives the adjoint system and the information wedge; Proposition 3.6 then decomposes equilibrium cost into a certainty-equivalent term and an information cost linked to endogenous precision.
@@
Recall the noise-state $\widehat{W}_{t}^{i}(\cdot)$ (2.5). We record the deterministic-kernel representation of the estimated-noise increment and its innovations gain in noise-state coordinates. Two-time kernels throughout live on the causal triangle $\Delta_{T}:=\{(t,s)\in[0,T]^{2}:0\leq s\leq t\leq T\}$, extended by 0 off $\Delta_{T}$. We write $d_{u}$ for increments in the path index $u$ at fixed estimation time $t$, and $d_{t}$ for updates in $t$ at fixed $u$. The state is assumed to admit the Volterra form $X_{t}=\bar{X}(t)+\int_{0}^{t}X(t,s)\,dW_{s}$, where $X(t,s)\in\mathbb{R}^{d\times(n+1)d}$ is a deterministic impulse-response kernel mapping primitive shocks into the physical state.
Theorem 3.1 (Volterra Filtering Closure).
Fix a player $i$. Assume the state admits the primitive-noise Volterra form
$X_{t}=\bar{X}(t)+\int_{0}^{t}X(t,s)\,dW_{s},$
@@
$F_{t}^{i}(u,s)=\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr.$ (3.2)
@@
$d_{t}\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}(t,u)^{\top}\sqrt{P^{i}(t)}\,dI_{t}^{i}\,du.$
@@
When opponents use noise-state Volterra strategies, a player’s best response over the full admissible $L^{2}$ class remains noise-state Volterra; the maximum principle yields a closed backward system of deterministic adjoint kernels including the information wedge $\mathcal{V}^{i}(t)$ that prices bilateral belief manipulation.
@@
Fix a player $i$ and suppose each opponent $k\neq i$ uses a noise-state Volterra strategy with deterministic kernels, which pins down a deterministic forward kernel environment (state impulse response $X(\cdot,\cdot)$, filtering objects $(P^{k},\widetilde{X}^{k},F^{k})$, and policy kernels $D^{k}$).
Structure. Because the equilibrium control is a Volterra process $D_{t}^{i}=\bar{D}^{i}(t)+\int_{0}^{t}D^{i}(t,u)\,dW_{u}$ with deterministic mean and kernel, the backward system splits into two structurally identical subsystems—one for the mean coefficients $(\bar{H}_{X}^{i},\bar{H}^{k,i})$, one for the kernel coefficients $(H_{X}^{i},H^{k,i})$—that should be read in parallel.
@@
Mean backward system. The mean physical costate $\bar{H}_{X}^{i}(t)$ and mean belief-adjoints $\bar{H}^{k,i}(t,u)$ satisfy
$\frac{d}{dt}\bar{H}_{X}^{i}(t)$ $=-\big(G^{i}_{XX}(t)\bar{X}(t)+G_{X}(t)\big)-A(t)^{\top}\bar{H}_{X}^{i}(t)-\underbrace{\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\widetilde{X}^{k}(t,z)\,\bar{H}^{k,i}(t,z)\,dz}_{=:\;\bar{\mathcal{V}}^{i}(t),\;\text{mean information wedge}},$ (3.3)
$\frac{d}{dt}\bar{H}^{k,i}(t,u)$ $=-\big(D^{k}(t,u)\big)^{\top}\bar{H}_{X}^{i}(t)+\int_{0}^{t}X(t,u)^{\top}P^{k}(t)\widetilde{X}^{k}(t,z)\,\bar{H}^{k,i}(t,z)\,dz,\qquad k\neq i,$ (3.4)
with terminal conditions $\bar{H}_{X}^{i}(T)=G^{i}_{XX}(T)\bar{X}(T)+G^{i}_{X}(T)$ and $\bar{H}^{k,i}(T,\cdot)=0$.
The first two terms in (3.3) are the standard single-agent costate equation; the third term $\bar{\mathcal{V}}^{i}(t)$ is the mean information wedge, absent in the single-agent problem. In the belief-adjoint equation (3.4), the forcing $D^{k}(t,u)^{\top}\bar{H}_{X}^{i}(t)$ transmits the physical costate into the belief channel through opponent $k$’s policy kernel: the larger $D^{k}(t,u)$, the more opponent $k$’s action at time $t$ responds to the $u$-th shock, and the more valuable it is for player $i$ to have shifted $k$’s estimate of that shock. The integral term is a self-coupling through which belief perturbations propagate forward via the filtering gain $\widetilde{X}^{k}$.
Kernel backward system. The kernel physical costate $H_{X}^{i}(t,r)$ and kernel belief-adjoints $H^{k,i}(t,u,r)$ satisfy the structurally identical system with $\bar{X}(t)\to X(t,r)$, $G_{X}(t)\to 0$, and $\bar{H}^{k,i}(t,u)\to H^{k,i}(t,u,r)$:
$\frac{d}{dt}H_{X}^{i}(t,r)$ $=-\,G^{i}_{XX}(t)X(t,r)-A(t)^{\top}H_{X}^{i}(t,r)-\underbrace{\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\widetilde{X}^{k}(t,z)\,H^{k,i}(t,z,r)\,dz}_{=:\;\mathcal{V}^{i}(t,r),\;\text{kernel information wedge}},$ (3.5)
$\frac{d}{dt}H^{k,i}(t,u,r)$ $=-\big(D^{k}(t,u)\big)^{\top}H_{X}^{i}(t,r)+\int_{0}^{t}X(t,u)^{\top}P^{k}(t)\widetilde{X}^{k}(t,z)\,H^{k,i}(t,z,r)\,dz,\qquad k\neq i,$ (3.6)
with terminal conditions $H_{X}^{i}(T,r)=G^{i}_{XX}(T)X(T,r)$ and $H^{k,i}(T,\cdot,r)=0$.
The two subsystems are identical in structure because they are the Volterra mean and kernel of a single stochastic costate. We define the (full) information wedge as the Volterra process
$\mathcal{V}^{i}_{t}:=\bar{\mathcal{V}}^{i}(t)+\int_{0}^{t}\mathcal{V}^{i}(t,r)\,dW_{r},$ (3.7)
which is the total distortion of the physical costate relative to the single-agent benchmark. The mean wedge $\bar{\mathcal{V}}^{i}(t)$ captures how belief manipulation distorts the equilibrium level (mean actions); the kernel wedge $\mathcal{V}^{i}(t,r)$ captures how it distorts the equilibrium response to shocks. Both components—and hence $\mathcal{V}^{i}_{t}$ itself—vanish when signals are exogenous.

### 18 Mar 2026 v2

@@ Abstract
In strategic environments with private information, evaluating a change in policy requires predicting how the equilibrium responds—but when actions reshape opponents’ signals, each agent’s optimal response depends on an infinite hierarchy of beliefs about beliefs [28] that has resisted exact analysis for four decades. We provide the first exact equilibrium characterization of finite-player continuous-time linear quadratic games with endogenous signals. Conditioning on primitive Brownian shocks rather than the physical state—a dynamic analogue of Harsanyi’s common-prior construction—collapses the belief hierarchy onto deterministic impulse response maps, reducing Nash equilibrium to a deterministic fixed point with no truncation and no large-population limit. The characterization yields an explicit information wedge that prices the marginal value of shifting opponents’ posteriors. The wedge vanishes precisely when signals are exogenous to controls, formally delineating the boundary where strategic belief manipulation matters, and provides a closed-form mapping from information primitives to equilibrium outcomes.
@@ 1 Introduction
This joint determination is the central obstacle. Consider a central bank shifting from quarterly to real-time disclosure of its output gap estimates. Each firm’s pricing decision depends on its belief about the gap and on its belief about how other firms filter the bank’s announcements. Under the new policy, signal precision changes; firms filter differently, revise their forecasts of one another’s pricing, and equilibrium inflation dynamics shift in tandem. Existing tools cannot trace this chain because the Townsend hierarchy makes the mapping from signal precision to equilibrium outcomes intractable. The characterization in this paper makes it explicit.
The same obstruction arises across a wide range of fields. In market microstructure, prices aggregate dispersed information and trading itself is informative. Dynamic information design faces the same difficulty: the mapping from signal structures to equilibrium outcomes becomes intractable once filtering and equilibrium interact. Contract theory encounters it when multiple agents exert unobservable effort on a shared outcome and draw inferences from it. So does decentralized control of autonomous vehicle fleets and multi-robot systems with misaligned operators, wherever agents observe a shared state through private channels and actions reshape what others sense. Existing approaches to the decentralized control problem typically assume cooperative teams or a common information base; the framework here requires neither.
@@
The conceptual move parallels Harsanyi’s [10] resolution of the belief hierarchy in incomplete-information games. Harsanyi replaced an infinite regress of beliefs about beliefs with a common prior over a type space from which all higher-order beliefs follow by conditioning. We replace the Townsend regress with the common probability space of primitive shocks $W=(W^{0},\ldots,W^{n})$, from which all higher-order conditional expectations follow by linear projection onto deterministic impulse response maps with no truncation. The analogy is imperfect in a revealing way: Harsanyi’s types are drawn once and the information partition is fixed, whereas here noise unfolds continuously and actions reshape signals, so the closure requires not just a common prior but a fixed point—the maps governing how projections evolve must be consistent with the strategy profile they induce (Figure 1).
@@
The main results establish that the best response amongst all admissible actions to any profile of linear strategies is itself linear. The equilibrium characterization produces an explicit information wedge $\mathcal{V}^{i}_{t}$, which prices the marginal value to player $i$ of shifting opponents’ posteriors. It vanishes identically when signals are exogenous to controls, formally delineating the boundary where strategic belief manipulation matters. The propagation $\Delta P^{i}\to\Delta\widetilde{X}^{i}\to\Delta\mathcal{V}^{i}_{t}\to\Delta D^{i}$ provides an explicit mapping from information primitives to equilibrium outcomes that a model with fixed filtering cannot deliver: because information enters through precision paths and affects the equilibrium only through deterministic impulse response maps, a principal can trace how disclosure precision propagates into equilibrium payoffs without simulating higher-order beliefs.
(i) Candidate Opponents use Noise-State Linear controls (ii) Closure State & filters have det. response maps (iii) Deviations Fixed maps $\Rightarrow$ det. propagation (iv) Best response FOC $\Rightarrow$ control noise-state linear (v) Fixed point BR map closes $\Longleftrightarrow$ Nash eq.
Figure 1: The Noise-State Linear fixed-point loop. (i) Assume opponents use Noise-state Linear controls. (ii) The state and filtering equations have deterministic impulse responses. (iii) Under fixed strategy maps, unobserved deviations propagate deterministically. (iv) The maximum principle produces a best response that is itself noise-state linear, closing the loop. (v) A fixed point of this best-response map is a Nash equilibrium.
@@ 1.1 Related Literature
The decentralized decision problem with shared objectives originates with Marschak and Radner [19]; Radner [23] established certainty equivalence and linear optimal policies for linear quadratic gaussian (LQG) teams. Team-theoretic solutions bypass the belief hierarchy because a social planner with aligned objectives solves a single-agent control problem—the fixed-point structure of Section 3 never arises. The information wedge $\mathcal{V}^{i}_{t}$ is an equilibrium object: it prices the value of shifting an opponent’s posterior, which requires distinct objectives to be well-defined.
@@
Nimark [22] approximates the higher-order belief hierarchy by truncation at finite order—the closest existing method in spirit to the present paper. The deterministic impulse-response system derived here provides an exact benchmark against which truncation error can be assessed.
@@ 2 The Decentralized LQG Game
The class of games we study has three features: (i) multiple agents share a common evolving state but observe it through private noisy channels; (ii) each agent’s action feeds back into the state dynamics and thereby into every other agent’s signal process; and (iii) agents minimize individual quadratic objectives over a finite horizon. Because actions enter the state, they simultaneously serve the physical objective and reshape what opponents observe, making each agent’s problem jointly strategic and informational. Features (i)–(ii) generate the infinite belief hierarchy identified by Townsend; feature (iii) provides the linear–Gaussian structure that makes it tractable—the same structure that underlies the workhorse models of [28, 26, 20] and the mean-field LQG literature [7], and that serves as the exact benchmark against which approximations can be assessed.
@@ 2.1 Primitives
Notation for primitive shocks.
@@
Unresolved impulse response.
@@
$\widetilde{X}^{i}_{t}(u):=X_{t}(u)-\widehat{X}^{i}_{t}(u),$
the component of the state’s impulse response to the $u$-th shock that player $i$ has not yet resolved at time tt—large when the shock is recent or the observation channel weak, and shrinking as observations accumulate. Formally, $\widetilde{X}^{i}_{t}(u)=\operatorname{Cov}(X_{t},\,dW_{u}\mid\mathcal{F}_{t}^{i})/du$; it serves as the filtering gain in the noise-state update (Theorem 3.1), with its algebraic representation in terms of $F^{i}$ given in (A.12).
@@
The linear state-cost coefficient $G^{X,i}_{t}=\bar{G}^{X,i}_{t}+\int_{0}^{t}G^{X,i}_{t}(r)\,dW_{r}$ is a stochastic process with deterministic mean $\bar{G}^{X,i}_{t}\in\mathbb{R}^{d}$ and deterministic kernel $G^{X,i}_{t}(r)\in\mathbb{R}^{d\times(n+1)d}$; deterministic coefficients are the special case $G^{X,i}_{t}(r)\equiv 0$. Each player $i$ incurs the random cost
$\begin{split}\mathcal{C}^{i}(D^{i};D^{-i}):=&\int_{0}^{T}\left(X_{t}^{\top}G^{XX,i}(t)X_{t}+2(G^{X,i}_{t})^{\top}X_{t}+G^{i}(t)+(D^{i}_{t})^{\top}G^{DD,i}(t)D^{i}_{t}\right)dt\\ &+X_{T}^{\top}G^{XX,i}(T)X_{T}+2(G^{X,i}_{T})^{\top}X_{T},\end{split}$ (2.3)
where $G^{XX,i}(t)$ and $G^{DD,i}(t)$ are symmetric positive semidefinite matrices (with $G^{DD,i}(t)$ positive definite), and $D^{-i}:=(D^{j})_{j\neq i}$ denotes opponents’ controls. Each player $i$ seeks to minimize
@@ 2.2 Nash Equilibrium
In a unilateral deviation by $i$, opponents’ strategy maps $Y^{k}_{\cdot\wedge t}\mapsto D_{t}^{k}[Y^{k}]$ are held fixed and applied to the perturbed observation histories; in the noise-state linear class these maps are parameterized by deterministic maps that remain unchanged off equilibrium.
@@
This paper aims to establish that the class of noise-state linear strategies is closed under best responses and to derive a deterministic impulse-response system whose fixed points correspond to Nash equilibria in that class.
Remark 2.2 (Subjective equilibrium and policy evaluation).
The best-response closure (Corollary 3.5) does not require player $i$ to know opponents’ maps correctly: a player holding incorrect beliefs about opponents’ strategies still solves a well-posed best-response problem against the believed profile. A policy-maker can therefore let each agent optimize under its own beliefs, then evaluate welfare through the realized state dynamics (2.1), which depend on the actual strategy profile rather than on whether agents’ beliefs about each other are correct. Common knowledge of rationality is sufficient for Nash equilibrium but is not required to apply the characterization as a best-response operator.
@@
Definition 2.2 ((Primitive-noise) Linear Process).
An $L^{2}$ stochastic process $L=(L_{t})_{t\in[0,T]}$ is a (primitive-noise) Linear process if there exist deterministic functions
@@
Remark 2.3 (Impulse-response notation).
We write $X_{t}(s)$ for the deterministic impulse response, not the random variable $X_{t}$ evaluated at an argument; the shock index $s$ distinguishes the two. The same subscript-$t$, parenthetical-shock convention applies to all two-time response maps.
Remark 2.4 (Properties of primitive-noise linear processes).
By construction, primitive-noise linear processes are $\mathbb{F}$-adapted Gaussian processes. Moreover, by Itô isometry the square-integrability condition above is equivalent to $L_{t}\in L^{2}$ for each $t$. Finally, the martingale representation theorem implies that any $L^{2}$ $\mathcal{F}_{t}$-measurable random variable admits a representation of the form $\int_{0}^{t}Z_{s}\,dW_{s}$ with a stochastic integrand $Z$; the defining restriction of the linear class is that the response map $L(t,\cdot)$ is deterministic (for each fixed $t$).
Definition 2.3 (Noise-state Linear control / strategy).
Fix a player $i$ and recall their noise-state path $\widehat{W}_{t}^{i}(\cdot)$ (Section 3). An admissible control $D^{i}=(D_{t}^{i})_{t\in[0,T]}$ is a noise-state linear control if it is affine in the noise state. That is, if there exist deterministic functions
@@
$D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{i}(u),\qquad\int_{0}^{t}\|D^{i}_{t}(u)\|^{2}\,du<\infty,$
and we extend $D^{i}_{t}(u)$ by 0 for $u>t$. Equivalently, writing $\widehat{W}^{\,i}[Y^{i}]$ to emphasize dependence on the observation path, this defines the strategy map
$Y^{i}_{\cdot\wedge t}\longmapsto D_{t}^{i}[Y^{i}]:=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{\,i}[Y^{i}](u).$
@@ 3 Main results
We work in the linear–Gaussian model of Section 2. Two results drive the analysis. First, when opponents use noise-state linear strategies, the filtering equations inherit deterministic impulse responses (Theorem 3.1). Second, the best response over the full admissible class is itself noise-state linear (Theorem 3.2), so Nash equilibrium reduces to a deterministic fixed point in impulse-response maps. The filtering closure (Theorem 3.1) shows that noise-state dynamics have deterministic kernels; the best-response closure (Theorem 3.2) derives the adjoint system and the information wedge; Proposition 3.6 then decomposes equilibrium cost into a certainty-equivalent term and an information cost linked to endogenous precision.
@@
Recall the noise-state $\widehat{W}_{t}^{i}(\cdot)$ (2.5). We record the deterministic-kernel representation of the estimated-noise increment and its innovations gain in noise-state coordinates. Two-time kernels throughout live on the causal triangle $\Delta_{T}:=\{(t,s)\in[0,T]^{2}:0\leq s\leq t\leq T\}$, extended by 0 off $\Delta_{T}$. We write $d_{u}$ for increments in the path index $u$ at fixed estimation time $t$, and $d_{t}$ for updates in $t$ at fixed $u$. The state is assumed to admit the linear form $X_{t}=\bar{X}_{t}+\int_{0}^{t}X_{t}(s)\,dW_{s}$, where $X_{t}(s)\in\mathbb{R}^{d\times(n+1)d}$ is a deterministic impulse-response kernel mapping primitive shocks into the physical state.
Theorem 3.1 (Linear Filtering Closure).
Fix a player $i$. Assume the state admits the primitive-noise linear form
$X_{t}=\bar{X}_{t}+\int_{0}^{t}X_{t}(s)\,dW_{s},$
@@
$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (3.2)
@@
$d_{t}\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,dI_{t}^{i}\,du.$
@@
When opponents use noise-state linear strategies, a player’s best response over the full admissible $L^{2}$ class remains noise-state linear; the maximum principle yields a closed backward system of deterministic adjoint kernels including the information wedge $\mathcal{V}^{i}(t)$ that prices bilateral belief manipulation.
@@
Fix a player $i$ and suppose each opponent $k\neq i$ uses a noise-state linear strategy with deterministic impulse-response maps, which pins down a deterministic forward environment (state impulse response $X(\cdot,\cdot)$, filtering objects $(P^{k},\widetilde{X}^{k},F^{k})$, and policy kernels $D^{k}$).
Structure. Because the equilibrium control is a noise-state linear process $D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,dW_{u}$ with deterministic mean and kernel, the backward system splits into two structurally identical subsystems—one for the mean coefficients $(\bar{H}^{X,i},\bar{H}^{k,i})$, one for the response-map coefficients $(H^{X,i},H^{k,i})$—that should be read in parallel.
@@
Mean backward system. The mean physical costate $\bar{H}^{X,i}_{t}$ and mean belief-adjoints $\bar{H}^{k,i}_{t}(u)$ satisfy
$\frac{d}{dt}\bar{H}^{X,i}_{t}$ $=-\big(G^{XX,i}(t)\bar{X}_{t}+\bar{G}^{X,i}_{t}\big)-A(t)^{\top}\bar{H}^{X,i}_{t}-\underbrace{\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz}_{=:\;\bar{\mathcal{V}}^{i}_{t},\;\text{mean information wedge}},$ (3.3)
$\frac{d}{dt}\bar{H}^{k,i}_{t}(u)$ $=-\big(D^{k}_{t}(u)\big)^{\top}\bar{H}^{X,i}_{t}+\int_{0}^{t}X_{t}(u)^{\top}P^{k}(t)\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz,\qquad k\neq i,$ (3.4)
with terminal conditions $\bar{H}^{X,i}_{T}=G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}$ and $\bar{H}^{k,i}_{T}(\cdot)=0$.
The first two terms in (3.3) are the standard single-agent costate equation; the third term $\bar{\mathcal{V}}^{i}_{t}$ is the mean information wedge, absent in the single-agent problem. In the belief-adjoint equation (3.4), the forcing $D^{k}_{t}(u)^{\top}\bar{H}^{X,i}_{t}$ transmits the physical costate into the belief channel through opponent $k$’s policy kernel: the larger $D^{k}_{t}(u)$, the more opponent $k$’s action at time $t$ responds to the $u$-th shock, and the more valuable it is for player $i$ to have shifted $k$’s estimate of that shock. The integral term is a self-coupling through which belief perturbations propagate forward via the filtering gain $\widetilde{X}^{k}$.
Kernel backward system. The kernel costate $H^{X,i}_{t}(r)$ and kernel belief-adjoints $H^{k,i}_{t}(u,r)$ satisfy the identical system with $\bar{X}_{t}\to X(t,r)$ and $\bar{G}^{X,i}_{t}\to G^{X,i}_{t}(r)$; the kernel information wedge $\mathcal{V}^{i}_{t}(r)$ is the corresponding substitution of $\bar{\mathcal{V}}^{i}_{t}$. We define the (full) information wedge as the linear process
$\mathcal{V}^{i}_{t}:=\bar{\mathcal{V}}^{i}_{t}+\int_{0}^{t}\mathcal{V}^{i}_{t}(r)\,dW_{r},$ (3.5)
which is the total distortion of the physical costate relative to the single-agent benchmark. The mean wedge $\bar{\mathcal{V}}^{i}_{t}$ captures how belief manipulation distorts the equilibrium level (mean actions); the response-map wedge $\mathcal{V}^{i}_{t}(r)$ captures how it distorts the equilibrium response to shocks. Both components—and hence $\mathcal{V}^{i}_{t}$ itself—vanish when signals are exogenous.